# Analyzing Multifiltering Functions Using Multiparameter Discrete Morse Theory

### v1

@@ 8. Discussion and closing remarks
Still, some questions remain open. From a theoretical perspective, as shown in smooth and continuous settings [5, 15, 21], the Pareto set of a vector filtering function is strongly related to the homological changes in the associated multifiltration. Thus, analogous connections could seemingly be made in our discrete setting. This could be addressed algebraically, by analyzing how the simplices in the Pareto set of a discrete multifiltering function are related in the associated multiparameter persistence modules [45], or dynamically, by considering a partition of level sets as a combinatorial multivector field [51]. Otherwise, by restraining the problem to the study of discrete vector functions induced by a component-wise injective vertex map, we could instead consider a piecewise linear approach, using the concepts defined in [28] or [37] as a starting point.
@@
Besides, there is room to improve some computational aspects of GenerateMDM. For instance, we could adapt the algorithm to multifiltering functions defined on a cubical complex (see [39]) so that it can be used to generate mdm functions on images. Also, we can verify that all mdm functions and gradient fields generated in Section 7, except those generated using $f=0$ as input, are perfect relatively to the input map. Hence, it would be of interest to find conditions on the input multifiltering function that guarantee the optimality of the output, as it was done in [60] for the algorithm ComputeDiscreteGradient. Besides, for GenerateMDM to be used to reduce the size of a simplicial complex prior to multipersistent homology computations, it should be improved in terms of speed. Indeed, using the multi-chunk algorithm as seen in [30, 32] or a minimal presentation of the homology of the complex [32, 41, 47] seems more efficient for this task.
@@
Nonetheless, GenerateMDM could potentially be used in other applications related to either multipersistent homology or topological data analysis. We could, for example, adapt the concepts of Morse-Smale complex [34], Morse connections graph [2] or descending complex [9] to mdm functions. Also, mdm functions and their gradient fields could realistically be used to modelize fluid dynamics, similarly to PL vector functions [37], or to extend the application of grayscale image analysis techniques [24, 36] to multispectral images.

### v3

@@ 7. Discussion and closing remarks
Still, some questions remain open. From a theoretical perspective, as shown in smooth and continuous settings [5, 15, 21], the Pareto set of a vector filtering function is strongly related to the homological changes in the associated multifiltration. Thus, analogous connections could seemingly be made in our discrete setting. This could be addressed algebraically, by analyzing how the simplices in the Pareto set of a discrete multifiltering function are related in the associated multiparameter persistence modules [46], or dynamically, by considering a partition into level sets as a combinatorial multivector field [52]. Otherwise, by restraining the problem to the study of discrete vector functions induced by a component-wise injective vertex map, we could instead consider a piecewise linear approach, using the concepts defined in [28] or [38] as a starting point.
@@
Furthermore, it would be possible to improve some computational aspects of GenerateMDM. For instance, we could adapt the algorithm to multifiltering functions defined on a cubical complex (see [40]) so that it can be used to generate mdm functions on images. Also, we may verify that almost every mdm function $g$ obtained experimentally in Section 6 is perfect relatively to its associated input map $f$, with only a few exceptions when $f=0$. Hence, it would be of interest to find theoretical conditions on the input of GenerateMDM that would guarantee the optimality of the output, as it was done in [46] for the algorithm ComputeDiscreteGradient (and thus for Matching, since it produces an equivalent output [63]). Besides, for GenerateMDM to be used to reduce the size of a simplicial complex prior to multipersistent homology computations, it should be improved in terms of speed. Indeed, using the multi-chunk algorithm as seen in [30, 32] or a minimal presentation of the homology of the complex [32, 42, 48] seems more efficient for this task.
@@
Nonetheless, GenerateMDM could potentially be used in other applications related to either multipersistent homology or topological data analysis. We could, for example, adapt the concepts of Morse-Smale complex [35], Morse connections graph [2] or descending complex [9] to mdm functions. Also, mdm functions and their gradient fields could realistically be used to modelize fluid dynamics, similarly to PL vector functions [38], or to extend the application of grayscale image analysis techniques [24, 37] to multispectral images. Finally, since multipersistent homology may be used to define topological descriptors of datasets in artificial intelligence applications [53], it would be of interest to see if mdm theory could also yield such descriptors.
Appendix A Proofs of Section 4.2
The main arguments which explain how the algorithms from Section 4.1 work were given in Section 4.2. Here, we present the comprehensive proofs of the correctness of those algorithms.
First, we show that every $\sigma\in K$ is processed exactly once. From Algorithm 1, it is easy to see that each level set $L_{u}$ is used exactly once as a parameter in ExpandMDM. Thus, it suffices to show that, for a given $L_{u}$, each $\sigma\in L_{u}$ is processed exactly once by ExpandMDM.
Lemma A.1.
Proof.
First, suppose $\sigma$ was processed as critical at line 26 of Algorithm 2. Then, $\sigma$ was in PQzero at some point, so all its facets have been processed and we have $\texttt{processed}(\alpha)=\textbf{True}$. In the case where $\sigma$ is paired with a facet $\tau<\sigma$, then $\sigma$ was popped from PQone at line 12 and $\tau=\texttt{unprocessed\_facet}(\sigma,L_{u})$. For all $\alpha\neq\tau$, we have $\texttt{processed}(\alpha)=\textbf{True}$ because $\tau$ is the unique unprocessed facet of $\sigma$. For $\alpha=\tau$, then $\alpha$ is paired with $\sigma$ which are both declared as processed simultaneously at line 19.
[…]
Now, assume $\sigma$ is the $n^{\text{th}}$ simplex in $L_{u}$ to be paired with a cofacet $\sigma^{\prime}$ and suppose that the first $n-1$ such simplices had no unprocessed facet in $L_{u}$ when they were paired. Consider a facet $\alpha<\sigma$ in $L_{u}$. As above, we could show that there exists a processed facet $\gamma\in L_{u}$ of $\sigma^{\prime}$ such that $\gamma\neq\sigma$ and $\alpha<\gamma<\sigma^{\prime}$.
Lemma A.2.
Proof.
[…]
From these observations, we can summarize the process of a simplex $\sigma\in L_{u}$ with the flowchart in Figure 17. Although it contains a few loops, we can show $\sigma$ always ends up being processed.
See 4.2
Proof.
The result follows from the three following statements:
  1. $\mathrm{(1)}$
    Each simplex in $L_{u}$ eventually enters PQone or PQzero.
  2. $\mathrm{(2)}$
    Every simplex in PQone and PQzero is eventually processed.
  3. $\mathrm{(3)}$
    A processed simplex cannot enter PQone or PQzero again.
  • At lines 7 and 9, all simplices of $L_{u}$ are still unprocessed.
  • At lines 20 and 27, a simplex $\sigma$ can only be added to PQone by function add_cofacets after one of its facet is processed. However, when $\texttt{processed}(\sigma)=\textbf{True}$, we know from Lemma A.1 that all facets of $\sigma$ are processed, so add_cofacets cannot add $\sigma$ to PQone.
    • if $\sigma$ does not have priority in PQone, another simplex is popped from PQone is either moved to PQzero or paired with a facet $\tau\in\texttt{PQzero}$ by Lemma A.2, so $\sigma$ remains unprocessed.
  • $L_{u}\subset K$ is finite;
  • simplices in PQzero cannot enter PQone;
  • processed simplices cannot return to either PQone or PQzero;
[…]
We now show that the dictionaries $g$ and ${\mathcal{V}}$ produced by GenerateMDM indeed represent a mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ and its gradient field ${\mathcal{V}}:K\nrightarrow K$. To do so, a few preliminary observations are needed.
Lemma A.3.
Proof.
Consider $u\in f(K)$ such that $\sigma\in L_{u}$. For a facet $\tau<\sigma$, we know that $f(\tau)\preceq f(\sigma)=u$, so either $f(\tau)=u^{\prime}\precneqq u$ or $f(\tau)=u$.
  • If $f(\tau)=u$, $\tau$ was processed within the same level set $L_{u}$ as $\sigma$, and we know from Lemma A.1 that $\tau$ cannot be processed before $\sigma$. ∎
Lemma A.4.
Proof.
Consider $u\in f(K)$ such that $\sigma_{0}\in L_{u}$. Since $\sigma_{0}$ is the first simplex to be processed, it belongs to the first level set $L_{u}$ given by the LevelSets routine, which implies that there exists no level set $L_{u^{\prime}}$ with $u^{\prime}\precneqq u$. Now, suppose $\sigma_{0}$ is processed as either one of the simplices $\sigma$ or $\tau$ from lines 1620. From the properties of a simplicial complex, we know $\sigma$ has at least another facet $\alpha\neq\tau$ and, since $f(\alpha)\preceq f(\sigma)=u$ and there exists no $u^{\prime}\precneqq u$, it follows that $\alpha\in L_{u}$. However, $\tau$ is the unique unprocessed facet of $\sigma$ in $L_{u}$, so $\texttt{processed}(\alpha)=\textbf{True}$, which is a contradiction since $\sigma_{0}\neq\alpha$ and $\sigma_{0}$ is the first simplex to be processed.
See 4.2
Proof.
For $n=0$, consider $u\in f(K)$ such that $\sigma_{0}\in L_{u}$. We know from Lemma A.4 that $\sigma_{0}$ is a critical vertex, so $g(\sigma_{0})$ is defined by calling $\texttt{ComputeG}(f,g,\delta,\sigma_{0})$ at line 24 of Algorithm 2. From Algorithm 3, it is obvious that $\texttt{ComputeG}(f,g,\delta,\sigma_{0})$ returns $f(\sigma_{0})\in{\mathbb{R}}^{k}$, so $g(\sigma_{0}):=f(\sigma_{0})$, which proves both $\mathrm{(1)}$ and $\mathrm{(3)}$. Also, because $\sigma_{0}$ is a vertex, it has no facet, which makes $\mathrm{(2)}$ trivial.
Now, consider $n\geq 1$ and suppose that the proposition is verified for simplices $\sigma_{0},\sigma_{1},...,\sigma_{n-1}$. If $\sigma_{n}$ is a critical vertex, we prove the proposition as above in the case $n=0$. Also, when $\sigma_{n}$ is processed as critical but not a vertex, the proof is similar to the case when $\sigma_{n}$ is paired with a facet by ExpandMDM. Thus, for the remainder of the proof, we assume $\sigma_{n}$ is processed as part of a pair $\tau<\sigma$. In this case, note that $g(\sigma_{n})$ is defined by $\texttt{ComputeG}(f,g,\delta,\sigma,\tau)$ at line 17 of Algorithm 2 and $\sigma$ is not a vertex, so ComputeG necessarily enters the else statement at line 5.
  • $\mathrm{(1)}$
    Consider the set $A$ of facets $\alpha<\sigma$ such that $\alpha\neq\tau$. Every $\alpha\in A$ was processed before $\sigma$ by Lemma A.3, so $g(\alpha)$ is well defined for each $\alpha\in A$ by the induction hypothesis. Hence, the instruction at line 7 and the if statement at line 10 of Algorithm 3 are executed without error, which proves $\mathrm{(1)}$.
  • $\mathrm{(2)}$
    • We first assume $\sigma_{n}=\sigma$. Consider the set $A$ of facets $\alpha<\sigma$ such that $\alpha\neq\tau$ and the vector value $w=(w_{1},...,w_{k})\in{\mathbb{R}}^{k}$ as defined at line 9 of Algorithm 3, so $w_{1}=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$ and $w_{i}=f_{i}(\sigma)$ for $i>1$. For $i=1$, it is clear that $w_{1}\geq g_{1}(\alpha)$ for all $\alpha\in A$, while for $i>1$,
      $w_{i}=f_{i}(\sigma)\geq f_{i}(\alpha)=g_{i}(\alpha)\ \forall\alpha\in A$
      because $f(\sigma)\succeq f(\alpha)$ by definition of an admissible map and we know that $f_{i}(\alpha)=g_{i}(\alpha)$ for all $i>1$ and all $\alpha\in A$ by the induction hypothesis. Thus, $w\succeq g(\alpha)$ for all $\alpha\in A$. Since $g_{1}(\sigma)$ is defined either as $w_{1}$ or $w_{1}+\delta$ and $g_{i}(\sigma):=w_{i}$ for $i>1$, it follows that $g(\sigma)\succeq w\succeq g(\alpha)$ for all $\alpha\in A$. Furthermore, if $w\succneqq g(\alpha)$ for all $\alpha\in A$, then $g(\sigma):=w\succneqq g(\alpha)$ for all $\alpha\in A$, and if $w=g(\alpha)$ for some $\alpha\in A$, then $g_{1}(\sigma):=w_{1}+\delta>w_{1}\geq g_{1}(\alpha)$. In both cases, we see that $g(\sigma)\succneqq g(\alpha)$ for all $\alpha\in A$, so if $g(\sigma)=g(\alpha)$ for some facet $\alpha<\sigma$, we necessarily have $\alpha\notin A$, and thus $\alpha=\tau$.
      Recall from line 17 of Algorithm 2 that $g(\tau)=g(\sigma)$. Moreover, we know from the properties of a simplicial complex that there exists a simplex $\alpha\neq\tau$ such that $\gamma<\alpha<\sigma$. Using the arguments from the previous case, we can see that $g(\sigma)\succneqq g(\alpha)$ since $\alpha$ is a facet of $\sigma$ such that $\alpha\neq\tau$. Also, $\alpha$ was processed before $\sigma$ and $\tau$ from Lemma A.3, so we can deduce from the induction hypothesis that the facet $\gamma$ of $\alpha$ is such that $g(\alpha)\succeq g(\gamma)$. We conclude that
      $g(\tau)=g(\sigma)\succneqq g(\alpha)\succeq g(\gamma).$
  • $\mathrm{(3)}$
    Recall that $g_{1}(\sigma_{n})$ is either $w_{1}$ or $w_{1}+\delta$ where $w_{1}$ is defined at line 7 of Algorithm 3 as $w_{1}=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$ and $A$ is the set of facets $\alpha<\sigma$ such that $\alpha\neq\tau$. Also, we see that $f_{1}(\sigma)=f_{1}(\tau)=f_{1}(\sigma_{n})$ since $\sigma_{n}$ is either $\sigma$ or $\tau$ and they both belong in a same level set.
    • Assume $w_{1}=f_{1}(\sigma)$. Then, we have either $g_{1}(\sigma_{n}):=f_{1}(\sigma_{n})$ or $g_{1}(\sigma_{n}):=f_{1}(\sigma_{n})+\delta$, so $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+m\delta$ where $m\in\{0,1\}$. Since $n\geq 1$ by hypothesis, we have $m\leq n$.
    • Assume $w_{1}=g_{1}(\alpha)>f_{1}(\sigma)$ for some $\alpha\in A$. By the induction hypothesis, $g_{1}(\alpha)=f_{1}(\alpha)+m\delta$ where $m$ is bounded by the number of simplices processed before $\alpha$. Thus, $w_{1}=f_{1}(\alpha)+m\delta$ and $m\leq n-1$. Also, we know that $f_{1}(\alpha)\leq f_{1}(\sigma)$ because $f$ is an admissible map.
      • $\bullet$
        If $f_{1}(\alpha)=f_{1}(\sigma)$, since $f_{1}(\sigma)=f_{1}(\sigma_{n})$ and $g_{1}(\sigma_{n})$ equals either $w_{1}$ or $w_{1}+\delta$, we necessarily have $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+m\delta$ or $g_{1}(\sigma_{n})=f_{1}(\sigma_{n})+(m+1)\delta$ where both $m,m+1\leq n$, hence the result.
      • $\bullet$
        We show the inequality $f_{1}(\alpha)<f_{1}(\sigma)$ leads to a contradiction. Recall from line 7 of Algorithm 1 that $0<\delta\leq\frac{f_{1}(\sigma)-f_{1}(\alpha)}{|K|}$. Also, $m<n<|K|$, so $m\delta<|K|\delta\leq f_{1}(\sigma)-f_{1}(\alpha)$. Therefore, $g_{1}(\alpha)=f_{1}(\alpha)+m\delta<f_{1}(\sigma)$, which contradicts the assumption that $w_{1}=g_{1}(\alpha)>f_{1}(\sigma)$. ∎
We are now ready to prove the main theorem.
See 4.2
Proof.
since, as defined in Algorithm 1, the parameter $\delta$ is such that $\delta\leq\frac{\epsilon}{|K|}$. Next, we show $g$ is a mdm function and ${\mathcal{V}}$ is its gradient field. Let $\sigma\in K$.
See 4.2
Proof.
If $f=g$, since $g$ is mdm, we obviously have $f$ mdm as well. Thus, we suppose $f$ is mdm and show that $f(\sigma)=g(\sigma)$ for all $\sigma\in K$.
As in the proof of Proposition 4.2, we label as $\{\sigma_{0},\sigma_{1},...,\sigma_{n},...,\sigma_{N}\}$ the simplices of $K$ so that $n<n^{\prime}$ when $\sigma_{n}$ is processed before $\sigma_{n^{\prime}}$ by GenerateMDM and we prove $f(\sigma_{n})=g(\sigma_{n})$ by induction on $n$.
Since $\sigma_{0}$ is a vertex processed as critical by Lemma A.4, we have that $g(\sigma_{0})$ is defined as $f(\sigma_{0})$, as seen at line 4 of Algorithm 3. For $n\geq 1$, if $\sigma_{n}$ is a vertex and processed as critical, we use the same argument to prove $g(\sigma_{n})=f(\sigma_{n})$. Otherwise, $g(\sigma_{n})$ is defined by going through lines 511 of Algorithm 3 either by processing $\sigma_{n}$ as a critical simplex or as part of a pair $\tau<\sigma$.
  • If $\sigma_{n}$ is part of a pair $\tau<\sigma$, then for all $\alpha\in A=\{\text{facets }\alpha<\sigma\ |\ \alpha\neq\tau\}$, we know that $\alpha$ is processed before $\sigma_{n}$, so $g_{1}(\alpha)=f_{1}(\alpha)$ by induction hypothesis. Hence, $w_{1}$ is defined at line 7 as
    $w_{1}:=$ $\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right),$
    $=$ $\max\left(\{f_{1}(\sigma)\}\cup\{f_{1}(\alpha)\ |\ \alpha\in A\}\right)=f_{1}(\sigma)$
    because $f(\sigma)\succeq f(\alpha)$ by definition of an admissible map $f$. Therefore, at line 9, $w$ is defined as $f(\sigma)$.
    Moreover, since $f$ is mdm, we know that $\operatorname{card}\mathcal{T}_{f}(\sigma)\leq 1$ and, considering that $f(\tau)=f(\sigma)$, we deduce $\mathcal{T}_{f}(\sigma)=\{\tau\}$. By definition of a mdm function, we have that $f(\sigma)\succneqq f(\alpha)$ for all $\alpha\in A=\{\text{facets }\alpha<\sigma\ |\ \alpha\neq\tau\}$. Hence, ComputeG does not enter the if statement at line 10, so $w=f(\sigma)$ is returned to define $g(\sigma_{n})$. Because $\sigma_{n}$ is either $\sigma$ or $\tau$ and $g(\sigma_{n})=f(\sigma)=f(\tau)$, we have the result.
  • If $\sigma_{n}$ is processed as a critical simplex, most of the previous reasoning still holds. More precisely, we see that for all $\alpha\in A=\{\text{facets }\alpha<\sigma_{n}\}$, we have $g_{1}(\alpha)=f_{1}(\alpha)$ by induction hypothesis and $w_{1}$ is defined as $w_{1}:=f_{1}(\sigma_{n})$ meaning that, at line 9, $w:=f(\sigma_{n})$.
    Furthermore, we could show that $\sigma_{n}$ being processed as critical means that $\sigma_{n}$ is a critical point of the mdm function $f$, thus $\operatorname{card}\mathcal{T}_{f}(\sigma_{n})=0$. Consequently, by definition of a mdm function, we have $f(\sigma_{n})\succneqq f(\alpha)$ for all facets $\alpha\in A$ and, again, ComputeG does not enter the if statement at line 10. We conclude that $w=f(\sigma_{n})$ is returned to define $g(\sigma_{n})$, hence the result.∎